OpenStudy (anonymous):

Graph the following system and locate the solution. 5y – 7x = 10 3y + 4x = 6

OpenStudy (anonymous):

First change both into slope-intercept form 5y - 7x = 10 + 7x + 7x 5y = 7x + 10 5y/5 = 7x/5 + 10/5 y = 7/5x + 2 3y + 4x = 6 -4x -4x 3y = -4x + 6 3y/3 = -4x/3 + 6/3 y = -4/3x + 2

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Then just graph it and find the solution. To graph it look at pic.